OpenStudy (mary.rojas):

1+2+3+4+5+6+7+8+9+10+11+12+13=

OpenStudy (alienium):

90

OpenStudy (mimi_x3):

\[\frac{n(n+1)}{2} = \frac{13(13+1)}{2} \]

OpenStudy (mary.rojas):

OMG @Mimi_x3 BINGOOOO!!!!

OpenStudy (mimi_x3):

:)